For the hydrogen atom, the energy of the electron is defined by the factor En=13.58/n2eV. Here n is positive integer. The minimum quantity of energy which it can absorb in its primitive stage is:

Grade:8

1 Answers

Mayank Ranka
2 years ago
En​=n2−13.58​
In primitive stage atn=1,E1​=−13.58eV
Minimum energy in excited stage,
n=2,E2​=−3.395eV
Energy absorbed=−3.395+13.58=10.19eV